OpenStudy (anonymous):

3 square root of 63 divided by 6 square root of 28 in simplest form

OpenStudy (blank ):

It is "3/4" if this is what the problem looks like: http://www4c.wolframalpha.com/Calculate/MSP/MSP17521ag82acf3621f8fc00003e87392dh180b278?MSPStoreType=image/gif&s=20&w=49.&h=46.

OpenStudy (anonymous):

yes it looks like that I just don't know how to show work for it

OpenStudy (blank ):

Just a second...

OpenStudy (blank ):

Well, I use the calculator on the computer to solve the sqrt parts. I did sqrt(63) = 7.9372539331937717715048472609178 7.937 * 3 = 23.811 sqrt(28) = 5.2915026221291811810032315072785 5.2915 * 6 = 31.749 23.811 / 31.749 = 0.749 or 0.75 0.75 is the same thing as 3/4 Hope that helps.

OpenStudy (anonymous):

ahh yess!! thank yu:)

OpenStudy (blank ):

ǝɯoɔlǝʍ ǝɹ,noʎ :)

OpenStudy (anonymous):

Refer to the attachment. Sorry. Sigma equation does not function on an IMac with the Chrome browser.

OpenStudy (blank ):

Thanks @robtobey

OpenStudy (anonymous):

You are welcome.

OpenStudy (anonymous):

thanks guys